OpenStudy (anonymous):

The dimensions of a rectangle are such that its length is 3 inches more than its width. If the length were doubled and if the width were decreased by 1 inch, the area would be decreased by 66 inches. What are the length and width of the rectangle.

OpenStudy (anonymous):

3 inches longer or shorter than the width?

OpenStudy (anonymous):

3inches more than width

OpenStudy (anonymous):

Sorry for the time it's taking but I think I got it now

OpenStudy (anonymous):

No, problem. Im n and out of the house doing chores.

OpenStudy (anonymous):

L = Length W = Width A = Area L*W = A (W + 3)(W) = A W^2 + 3W = A 2(W + 3))*(W - 1) = 2(W^2 + 2W - 3) = 2W^2 + 4W - 6 The change in area is equal to the new Area minus the original Area (2W^2 + 4W - 6) - (W^2 + 3W) = 66 W^2 + W - 6 = 66 W^2 + W - 72 = 0 (W - 8)(W + 9) = 0 I.e width => W = 8 or W = -9 As Width cannot be negative W = 8 and L = 8 + 3 L = 11 I hope that works out right!

OpenStudy (anonymous):

Well, if not I really appreciate the time and effort that you put in But, I think that it is and I will let you know